Grade 12th passAlgebra

Find the square of the following numbers.
93..........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

(93)2
= (90+3 )2
= (90)2 + (3)2 + 2×90×3 [Since, (a+b)2 = a2+b2 +2ab]
=8100 + 9 + 540
= 8649

Thanks